# Task-agnostic Pre-training and Task-guided Fine-tuning for Versatile Diffusion Planner

### v1

@@ Abstract
Diffusion models have demonstrated their capabilities in modeling trajectories of multi-tasks. However, existing multi-task planners or policies typically rely on task-specific demonstrations via multi-task imitation, or require task-specific reward labels to facilitate policy optimization via Reinforcement Learning (RL). To address these challenges, we aim to develop a versatile diffusion planner that can leverage large-scale inferior data that contains task-agnostic sub-optimal trajectories, with the ability to fast adapt to specific tasks. In this paper, we propose SODP, a two-stage framework that leverages Sub-Optimal data to learn a Diffusion Planner, which is generalizable for various downstream tasks. Specifically, in the pre-training stage, we train a foundation diffusion planner that extracts general planning capabilities by modeling the versatile distribution of multi-task trajectories, which can be sub-optimal and has wide data coverage. Then for downstream tasks, we adopt RL-based fine-tuning with task-specific rewards to fast refine the diffusion planner, which aims to generate action sequences with higher task-specific returns. Experimental results from multi-task domains including Meta-World and Adroit demonstrate that SODP outperforms state-of-the-art methods with only a small amount of data for reward-guided fine-tuning.
@@ 1 Introduction
There has been a long-standing pursuit to develop agents capable of performing multiple tasks (Reed et al., 2022; Lee et al., 2022). While traditional RL methods have made significant strides in training agents to master individual tasks (Silver et al., 2016; OpenAI et al., 2019), extending this capability to handle diverse tasks remains a significant challenge due to the diversity of task variants and optimization directions with different rewards. Multi-task RL aims to address this by developing agents via task-conditioned optimization (Yu et al., 2020; Lee et al., 2022) or parameter-compositional learning (Sun et al., 2022; Lee et al., 2023). However, the inherent diversity in task trajectory distributions makes it challenging to model and accommodate the modeling across different task structures. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), originally designed for generative tasks, provide a powerful framework for addressing these difficulties. Their capacities to capture complex, multi-modal distributions within high-dimensional data spaces (Podell et al., 2023; Ho et al., 2022; Jing et al., 2022) makes them well-suited for representing the broad variability encountered in multi-task environments.
Motivated by this, existing methods have employed diffusion models to imitate expert behaviors derived from human demonstrations across various tasks (Pearce et al., 2023; Xu et al., 2023; Chi et al., 2023). However, acquiring task-specific demonstrations is often time-consuming and costly, especially in environments requiring specialized domain expertise. Alternative approaches attempt to optimize diffusion models with return guidance (He et al., 2024; Liang et al., 2023) or conventional RL paradigm (Wang et al., 2022b), which demands a large volume of data with reward labels for each task. To address the above limitations, we wonder whether a generalized diffusion planner can be learned from a large amount of low-quality trajectories without reward labels, with the ability to adapt quickly to various downstream tasks. The inferior dataset comprises a mixture of sub-optimal state-action pairs from various tasks, which can be easily obtained in the real world. In training, the diffusion planner seeks to model the distribution of diverse trajectories with broad coverage, enabling it to acquire generalizable capabilities and allowing the planner to further concentrate on high-reward regions of specific downstream tasks via fast adaptation. An overview of our method is given in Figure 1.
@@
In this paper, we propose a novel framework to utilize Sub-Optimal data to train a Diffusion Planner (SODP) that can generalize across a wide range of downstream tasks. SODP consists of two stages: pre-training and fine-tuning. By leveraging a set of trajectories of different tasks for pre-training, we employ action-sequence prediction to capture shared knowledge across tasks. Since the state space may vary between tasks, focusing on the common action space (e.g., end-effector poses of a robot arm) facilitates task generalization. We frame the pre-training stage as a conditional generative problem that generates future actions based on historical states. Then, inspired by the remarkable success of RL-based alignment for LLMs (Ouyang et al., 2022; Glaese et al., 2022), we adopt an RL-based fine-tuning approach to tailor the pre-trained diffusion planner to specific downstream tasks. Specifically, we conduct online interaction based on the pre-trained planner to collect task-specific experiences with reward labels, and perform policy gradients to iteratively refine the predicted action-sequence distribution based on reward feedback of tasks. Through fine-tuning, the diffusion planner can gradually adapt toward generating actions with high task-specific rewards and eventually become optimal for the given task.
@@
Figure 2 illustrates our method. In pre-training, the model captures diverse behavior patterns from the training data, encompassing inferior and superior actions. After fine-tuning, the model shrinks the action distribution and concentrates on generating optimal action sequences for a specific task. Our contributions can be summarized as follows. (i) We propose a novel pre-training and fine-tuning paradigm for learning a versatile diffusion planner, which leverages sub-optimal transitions to capture the broad action distributions across tasks, and adopt task-specific fine-tuning to transfer the planner to downstream tasks. (ii) We give an efficient fine-tuning algorithm based on policy gradient for diffusion planners, which progressively shifts the action distribution to concentrate on regions associated with higher task returns. (iii) We conduct extensive experiments using sub-optimal data from state-based Meta-World (Yu et al., 2019) as well as image-based Adroit (Rajeswaran et al., 2017), showcasing its superior performance compared to state-of-the-art approaches.
@@ 2 Preliminaries
$q(\bm{x}_{k}|\bm{x}_{k-1}):=\mathcal{N}(\bm{x}_{k};\sqrt{1-\beta_{k}}\bm{x}_{k-1},\beta_{k}\bm{I}),$ (1)
@@
$\mathcal{L}_{\text{denoise}}(\theta):=\mathbb{E}_{k\sim[1,K],x_{0}\sim q,\epsilon\sim\mathcal{N}(0,\bm{I})}\left[\big{\|}\epsilon-\epsilon_{\theta}(\bm{x}_{k},k)\big{\|}^{2}\right].$ (3)
@@
$\bm{x}_{k-1}=\frac{1}{\sqrt{\alpha_{k}}}\left(\bm{x}_{k}-\frac{1-\alpha_{k}}{\sqrt{1-\bar{\alpha}_{k}}}\epsilon_{\theta}(\bm{x}_{k},k)\right)+\sigma_{k}\mathcal{N}(0,\bm{I}),$ (4)
@@ 3.1 Pre-training with Large-scale Sub-optimal Data
Previous works (He et al., 2024) typically model multi-task RL as a conditional generative problem using diffusion models trained on datasets composed of multiple task subsets $\mathcal{D}=\cup_{i=1}^{N}\mathcal{D}_{i}$, as:
@@
Here, we represent $\bm{x}_{0}:=\bm{a}_{t}^{0}=(a_{t},a_{t+1},...,a_{t+H-1})$ as an action sequence, where $H$ is the planning horizon and $t$ is the timestep sampled from dataset $\mathcal{D}$. As previous work (Chi et al., 2023), we denote $\bm{s}_{t}$ as the historical states at timestep $t$ with length $T_{o}$, i.e., $\bm{s}_{t}:=\{s_{t-T_{o}+1},\dots s_{t-1},s_{t}\}$. The formulation in Eq. (6) enables the model to learn the broad action-sequence distribution of multi-tasks depending on previous observations, without requiring additional guidance. To train our planning model, we modify Eq. (3) to obtain our pre-training objective as follows:
$\mathcal{L}_{\text{pre-train}}(\theta)=\mathbb{E}_{k\sim[1,K],(\bm{s}_{t},\bm{a}^{0}_{t})\sim D,\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}(\bm{a}_{t}^{k},\bm{s}_{t},k)\right\|^{2}\right].$ (7)
@@
$\bm{a}^{k-1}_{t}=\frac{1}{\sqrt{\alpha_{k}}}\left(\bm{a}^{k}_{t}-\frac{1-\alpha_{k}}{\sqrt{1-\bar{\alpha}_{k}}}\epsilon_{\theta}(\bm{a}_{t}^{k},\bm{s}_{t},k)\right)+\sigma_{k}\mathcal{N}(0,\bm{I}).$ (8)
@@ 3.2 Reward Fine-tuning for Downstream Tasks
$s_{k}^{\text{diff}}=(s_{t},\bm{a}_{t}^{K-k}),\quad\;\;a_{k}^{\text{diff}}=\bm{a}_{t}^{K-k-1},\quad\;\;P_{0}^{\text{diff}}(s_{0}^{\text{diff}})=(\delta_{s_{t}},\mathcal{N}(0,I)),$
$P^{\text{diff}}(s_{k+1}^{\text{diff}}\mid s_{k}^{\text{diff}},a_{k}^{\text{diff}})=(\delta_{s_{t}},\delta_{a_{k}^{\text{diff}}}),\quad\;\;R^{\text{diff}}(s_{k}^{\text{diff}},a_{k}^{\text{diff}})=\begin{cases}r(s_{k+1}^{\text{diff}})=r(\bm{a}_{t}^{0})&\text{if }\;k=K-1,\\ 0&\text{otherwise}.\end{cases},$
@@
where $s_{k}^{\text{diff}}$ and $a_{k}^{\text{diff}}$ are the state and action at timestep $k$, $P_{0}^{\text{diff}}$ and $P^{\text{diff}}$ are the initial distribution and transition dynamics, $\delta$ is the Dirac delta distribution, $R^{\text{diff}}$ is the reward function and $p_{\theta}(\bm{a}_{t}^{K-k-1}\mid\bm{a}_{t}^{K-k},\bm{s}_{t})$ is the pre-trained diffusion planner. This formulation allows the state transitions in the MDP to be mapped to the denoising process in the diffusion model. The MDP initiates by sampling an initial state $s_{0}^{\text{diff}}\sim P_{0}^{\text{diff}}$, which corresponds to sample Gaussian noise $\bm{a}^{K}_{t}$ at the beginning of the reverse process. At each timestep $k$, the policy $\pi_{\theta}^{\text{diff}}(a_{k}^{\text{diff}}\mid s_{k}^{\text{diff}})$ takes an action $a_{k}^{\text{diff}}$ based on current state $s_{k}^{\text{diff}}$, which corresponds to generate next latent $\bm{a}_{t}^{K-k-1}$ based on current latent $\bm{a}_{t}^{K-k}$ following Eq. (8). The reward remains zero until a noise-free output $\bm{a_{0}^{t}}$ is evaluated. Different from previous text-to-image studies that typically evaluate the final sample using a pre-trained reward model (Black et al., 2023; Fan et al., 2024), we aim to fine-tune the pre-trained diffusion planner to maximize rewards of downstream tasks, which makes constructing reward models for all tasks costly. Therefore, we directly evaluate the generated action sequences in an online RL environment for each specific task $\mathcal{T}$. Specifically, for any given timestep $t$, we use the planner to generate future actions $\bm{a}_{t}^{0}=(a_{t},a_{t+1},\dots,a_{t+H-1})$ and then execute the first $T_{a}$ steps. Then we calculate the discounted cumulative reward from the environment to assess the generated sample, expressed as $r(\bm{a}_{t}^{0})=\sum\nolimits_{t}^{T_{a}}\gamma^{t-1}r^{\mathcal{T}}(s_{t},a_{t})$. We write $r(\bm{a}_{t}^{0})$ as shorthand for $r(s_{t},\bm{a}_{t}^{0})$ for brevity.
Finetuning objective.
@@
$\nabla_{\theta}J^{\mathcal{T}}(\theta)=\sum\nolimits_{t}\mathbb{E}_{p_{\theta}(\bm{a}_{t}^{0:K}|\bm{s}_{t})}\left[r^{\mathcal{T}}(\bm{a}^{0}_{t})\sum_{k=1}^{K}\nabla_{\theta}\log p_{\theta}(\bm{a}_{t}^{k-1}|\bm{a}_{t}^{k},\bm{s}_{t})\right].$ (11)
However, optimizing with Eq. (11) can be computationally intensive, as it requires generating new samples after each optimization step. To enhance sample efficiency and leverage historical sequences, we employ importance sampling, following the approach of proximal policy optimization (PPO) (Schulman et al., 2017), and derive the loss function for reward improvement as follows:
$\mathcal{L}_{\text{Imp}}^{\mathcal{T}}(\theta)=\sum\nolimits_{t}\mathbb{E}_{p_{\theta_{\text{old}}}(\bm{a}_{t}^{0:K}|\bm{s}_{t})}\left[\sum_{k=1}^{K}-r^{\mathcal{T}}(\bm{a}^{0}_{t})\text{max}\Big{(}\rho_{k}(\theta,\theta_{\text{old}}),\text{clip}\left(\rho_{k}(\theta,\theta_{\text{old}}),1+\epsilon,1-\epsilon\right)\Big{)}\right],$ (12)
@@
However, fine-tuning the model solely depending on the reward is insufficient since the model may step too far, which can lead to performance collapse and instability during reward maximization. To address this problem, we introduce a Behavior-Clone (BC) regularization term during the fine-tuning process. Concretely, we aim to constrain our policy $\theta$ to closely match a target policy $\mu$, ensuring that $\theta$ does not deviate significantly from $\mu$ after policy updates. This constraint can be modeled using a negative log-likelihood (NLL) loss as:
@@
where $\epsilon(\bm{a}_{\mu}^{k},k)$ represents the ground-truth noise added to $\bm{a}_{\mu}^{k}$ at timestep $k$, which can be calculated as $\epsilon(\bm{a}_{\mu}^{k},k)=\frac{\bm{a}_{\mu}^{k}-\sqrt{\bar{\alpha}_{k}}\cdot\bm{a}_{\mu}^{0}}{\sqrt{1-\bar{\alpha}_{k}}}.$
How to select the target policy? Intuitively, an ideal target policy is the optimal policy that generates samples $x^{*}$ satisfying $\mathcal{C}(\bm{x}^{*})\geq\mathcal{C}(\bm{x})$ for all possible $\bm{x}$, where $\mathcal{C}(\bm{x})$ represents a measure of the performance or quality of the sample, such as the accumulated reward for action sequences. Since $\mu$ is unknown during fine-tuning, we approximate it by sampling action sequences $\bm{a}$ that satisfy $\mathcal{C}(\bm{a})\approx\mathcal{C}(\bm{a}^{*})$. In practice, we denote $\bm{a}^{*}$ as the best actions from recent play experience, such as those that yielded the top $n$ highest rewards or successfully completed the given task. We then sample $\bm{a}^{k}$ from these proficient action sequences obtained during online interaction, nearly equivalent to sampling from $\mu$ to regularize the fine-tuning process. We also remark that the BC regularizer is not the only way to incorporate regularization into Eq. (12). For example, a Kullback–Leibler (KL) divergence between the fine-tuned and pre-trained models, or a diffusion pre-train loss can be employed to regularize the fine-tuning process, as shown in text-to-image and text-to-speech generation (Fan et al., 2024; Chen et al., 2024). However, we find these regularization may cause the pre-trained planner trap in sub-optimal regions, hindering performance improvement. We will further discuss them in experiments.
Combining Eq. (12) with Eq. (14), the loss function for reward fine-tuning in downstream tasks $\mathcal{T}\sim p(\mathcal{T})$ is expressed as follows:
@@
where $\lambda$ is a weight coefficient. The overall process of pre-training and fine-tuning using SODP is summarized in Alg. 1 in the appendix. Since our goal is to generate complete trajectories rather than individual segments, we utilize a trajectory-level buffer (Zheng et al., 2022) for estimating the target policy $\mu$. Further, to ensure the accuracy of the approximation, we generate several proficient trajectories using the pre-trained model at the beginning of each iteration.

### v3

@@ Abstract
Diffusion models have demonstrated their capabilities in modeling trajectories of multi-tasks. However, existing multi-task planners or policies typically rely on task-specific demonstrations via multi-task imitation, or require task-specific reward labels to facilitate policy optimization via Reinforcement Learning (RL). They are costly due to the substantial human efforts required to collect expert data or design reward functions. To address these challenges, we aim to develop a versatile diffusion planner capable of leveraging large-scale inferior data that contains task-agnostic sub-optimal trajectories, with the ability to fast adapt to specific tasks. In this paper, we propose SODP, a two-stage framework that leverages Sub-Optimal data to learn a Diffusion Planner, which is generalizable for various downstream tasks. Specifically, in the pre-training stage, we train a foundation diffusion planner that extracts general planning capabilities by modeling the versatile distribution of multi-task trajectories, which can be sub-optimal and has wide data coverage. Then for downstream tasks, we adopt RL-based fine-tuning with task-specific rewards to quickly refine the diffusion planner, which aims to generate action sequences with higher task-specific returns. Experimental results from multi-task domains including Meta-World and Adroit demonstrate that SODP outperforms state-of-the-art methods with only a small amount of data for reward-guided fine-tuning.
Machine Learning, ICML
@@ 1 Introduction
There has been a long-standing pursuit to develop agents capable of performing multiple tasks (Reed et al., 2022; Lee et al., 2022). Although traditional RL methods have made significant strides in training agents to master individual tasks (Silver et al., 2016; OpenAI et al., 2019), expanding this capability to handle diverse tasks remains a significant challenge due to the diversity of task variants and optimization directions with different rewards. Multi-task RL aims to address this by developing agents via task-conditioned optimization (Yu et al., 2020; Lee et al., 2022) or parameter-compositional learning (Sun et al., 2022; Lee et al., 2023). However, existing approaches often struggle with policy representation since they assume shared latent structures across tasks, which causes issues when tasks have significantly different dynamics and rewards, limiting their ability to capture the multi-modal distribution of optimal behaviors across diverse task spaces. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), originally designed for generative tasks, provide a powerful framework to address these difficulties. Their capacity to capture complex, multi-modal distributions within high-dimensional data spaces (Podell et al., 2024; Ho et al., 2022; Jing et al., 2022) makes them well suited to represent the broad variability encountered in multi-task environments.
@@
Motivated by this, existing methods have employed diffusion models to mimic expert behaviors derived from human demonstrations on various tasks (Pearce et al., 2023; Xu et al., 2023; Chi et al., 2023). However, acquiring task-specific demonstrations is often time-consuming and costly, especially in environments requiring specialized domain expertise. Alternative approaches attempt to optimize diffusion models with return guidance (He et al., 2023; Liang et al., 2023) or conventional RL paradigm (Wang et al., 2023), which demands a large volume of data with reward labels for each task. To address the above limitations, we wonder whether a generalized diffusion planner can be learned from a large amount of low-quality trajectories without reward labels, with the ability to adapt quickly to various downstream tasks. We only require the inferior data to comprise a mixture of sub-optimal state-action pairs from various tasks, which can be easily obtained in the real world. In training, the diffusion planner seeks to model the distribution of diverse trajectories with broad coverage, enabling it to acquire generalizable capabilities and allowing the planner to further concentrate on high-reward regions of specific downstream tasks via fast adaptation. An overview of our method is given in Figure 1.
In this paper, we propose a novel framework to utilize Sub-Optimal data to train a Diffusion Planner (SODP) that can generalize across a wide range of downstream tasks. SODP consists of two stages: pre-training and fine-tuning. By leveraging a set of trajectories of different tasks for pre-training, we employ action-sequence prediction to capture shared knowledge across tasks. Since the state space may vary between tasks, focusing on the common action space (e.g., end-effector poses of a robot arm) facilitates task generalization. We frame the pre-training stage as a conditional generative problem that generates future actions based on historical states. Then, inspired by the remarkable success of RL-based alignment for LLMs (Ouyang et al., 2022; Bai et al., 2025), we adopt an RL-based fine-tuning approach to tailor the pre-trained diffusion planner to specific downstream tasks. Specifically, we conduct online interaction based on the pre-trained planner to collect task-specific experiences with reward labels, and perform policy gradients to iteratively refine the predicted action-sequence distribution based on reward feedback of tasks. Through fine-tuning, the diffusion planner can gradually adapt toward generating actions with high task-specific rewards and eventually become optimal for the given task. Figure 2 illustrates our method. In pre-training, the model captures diverse behavior patterns from training data, encompassing inferior and mediocre actions. After fine-tuning, the model shrinks the action distribution and concentrates on generating optimal action sequences for a specific task. Our contributions can be summarized as follows. (i) We propose a novel pre-training and fine-tuning paradigm for learning a versatile diffusion planner, which leverages sub-optimal transitions to capture the broad action distributions across tasks, and adopt task-specific fine-tuning to transfer the planner to downstream tasks. (ii) We give an efficient fine-tuning algorithm based on policy gradient for diffusion planners, which progressively shifts the action distribution to concentrate on regions associated with higher task returns. (iii) We conduct extensive experiments using sub-optimal data from Meta-World (Yu et al., 2019), showcasing its superior performance compared to state-of-the-art approaches.
@@ 2 Preliminaries
$q(\bm{x}_{k}|\bm{x}_{k-1}):=\mathcal{N}(\bm{x}_{k};\sqrt{1-\beta_{k}}\bm{x}_{k-1},\beta_{k}\boldsymbol{I}),$ (1)
@@
$\mathcal{L}_{\text{denoise}}(\theta):=\mathbb{E}_{k\sim[1,K],x_{0}\sim q,\epsilon\sim\mathcal{N}(0,\boldsymbol{I})}\left[\big{\|}\epsilon-\epsilon_{\theta}(\bm{x}_{k},k)\big{\|}^{2}\right].$ (3)
@@
$\bm{x}_{k-1}=\frac{1}{\sqrt{\alpha_{k}}}\left(\bm{x}_{k}-\frac{1-\alpha_{k}}{\sqrt{1-\bar{\alpha}_{k}}}\epsilon_{\theta}(\bm{x}_{k},k)\right)+\sigma_{k}\mathcal{N}(0,\boldsymbol{I}),$ (4)
@@ 3.1 Pre-training with Large-scale Sub-optimal Data
Previous works typically model multi-task RL as a conditional generative problem using diffusion models trained on datasets composed of multiple task subsets $\mathcal{D}=\cup_{i=1}^{N}\mathcal{D}_{i}$, as:
@@
Here, we represent $\bm{x}_{0}:=\bm{a}_{t}^{0}=(a_{t},a_{t+1},...,a_{t+H-1})$ as an action sequence, where $H$ is the planning horizon and $t$ is the timestep sampled from dataset $\mathcal{D}$. We denote $\bm{s}_{t}$ as the historical states at timestep $t$ with length $T_{o}$, i.e., $\bm{s}_{t}:=\{s_{t-T_{o}+1},\dots s_{t-1},s_{t}\}$. The formulation in Eq. (6) enables the model to learn the broad action-sequence distribution of multi-tasks depending on previous observations, without requiring additional guidance. To train our planning model, we modify Eq. (3) to obtain our pre-training objective as:
$\mathcal{L}_{\text{pre-train}}(\theta)=\mathbb{E}_{k,\epsilon,(\bm{s}_{t},\bm{a}^{0}_{t})\sim D}\left[\left\|\epsilon-\epsilon_{\theta}(\bm{a}_{t}^{k},\bm{s}_{t},k)\right\|^{2}\right].$ (7)
@@
$\bm{a}^{k-1}_{t}=\frac{1}{\sqrt{\alpha_{k}}}\left(\bm{a}^{k}_{t}-\frac{1-\alpha_{k}}{\sqrt{1-\bar{\alpha}_{k}}}\epsilon_{\theta}(\bm{a}_{t}^{k},\bm{s}_{t},k)\right)+\sigma_{k}\mathcal{N}(0,\boldsymbol{I}).$ (8)
@@ 3.2 Reward Fine-tuning for Downstream Tasks
$s_{k}^{\text{diff}}=(s_{t},\bm{a}_{t}^{K-k}),\quad\;\;a_{k}^{\text{diff}}=\bm{a}_{t}^{K-k-1},$
$P_{0}^{\text{diff}}(s_{0}^{\text{diff}})=(\delta_{s_{t}},\mathcal{N}(0,I)),$
$P^{\text{diff}}(s_{k+1}^{\text{diff}}\mid s_{k}^{\text{diff}},a_{k}^{\text{diff}})=(\delta_{s_{t}},\delta_{a_{k}^{\text{diff}}}),$
$R^{\text{diff}}(s_{k}^{\text{diff}},a_{k}^{\text{diff}})=\begin{cases}r(s_{k+1}^{\text{diff}})=r(\bm{a}_{t}^{0})&\text{if }\;k=K-1,\\ 0&\text{otherwise}.\end{cases},$
@@
where $s_{k}^{\text{diff}}$ and $a_{k}^{\text{diff}}$ are the state and action at timestep $k$, $P_{0}^{\text{diff}}$ and $P^{\text{diff}}$ are the initial distribution and transition dynamics, $\delta$ is the Dirac delta distribution, $R^{\text{diff}}$ is the reward function and $p_{\theta}(\bm{a}_{t}^{K-k-1}\mid\bm{a}_{t}^{K-k},\bm{s}_{t})$ is the pre-trained diffusion planner. This formulation allows the state transitions in the MDP to be mapped to the denoising process in the diffusion model. The MDP initiates by sampling an initial state $s_{0}^{\text{diff}}\sim P_{0}^{\text{diff}}$, which corresponds to sample Gaussian noise $\bm{a}^{K}_{t}$ at the beginning of the reverse process. At each timestep $k$, the policy $\pi_{\theta}^{\text{diff}}(a_{k}^{\text{diff}}\mid s_{k}^{\text{diff}})$ takes an action $a_{k}^{\text{diff}}$ based on current state $s_{k}^{\text{diff}}$, which corresponds to generate next latent $\bm{a}_{t}^{K-k-1}$ based on current latent $\bm{a}_{t}^{K-k}$ following Eq. (8). The reward remains zero until a noise-free output $\bm{a^{0}_{t}}$ is evaluated. Different from previous text-to-image studies that typically evaluate the final sample using a pre-trained reward model (Black et al., 2024; Fan et al., 2024), we aim to fine-tune the pre-trained diffusion planner to maximize rewards of downstream tasks, which makes constructing reward models for all tasks costly. Therefore, we directly evaluate the generated action sequences in an online RL environment for each specific task $\mathcal{T}$. Specifically, for any given timestep $t$, we use the planner to generate future actions $\bm{a}_{t}^{0}=(a_{t},a_{t+1},\dots,a_{t+H-1})$ and then execute the first $T_{a}$ steps. Then we calculate the discounted cumulative reward from the environment to assess the generated sample, expressed as $r(\bm{a}_{t}^{0})=\sum\nolimits_{t}^{T_{a}}\gamma^{t-1}r^{\mathcal{T}}(s_{t},a_{t})$. We write $r(\bm{a}_{t}^{0})$ as shorthand for $r(s_{t},\bm{a}_{t}^{0})$ for brevity.
Fine-tuning objective.
@@
$\nabla_{\theta}J^{\mathcal{T}}(\theta)=$
$\sum\nolimits_{t}\mathbb{E}_{p_{\theta}(\bm{a}_{t}^{0:K}|\bm{s}_{t})}\left[r^{\mathcal{T}}(\bm{a}^{0}_{t})\sum_{k=1}^{K}\nabla_{\theta}\log p_{\theta}(\bm{a}_{t}^{k-1}|\bm{a}_{t}^{k},\bm{s}_{t})\right].$ (11)
However, optimizing with Eq. (3.2) can be computationally intensive, as it requires generating new samples after each optimization step. To enhance sample efficiency and leverage historical sequences, we employ importance sampling, following the approach of proximal policy optimization (PPO) (Schulman et al., 2017), and derive the loss function for reward improvement as follows:
$\mathcal{L}_{\text{Imp}}^{\mathcal{T}}(\theta)=\sum\nolimits_{t}\mathbb{E}_{p_{\theta_{\text{old}}}(\bm{a}_{t}^{0:K}|\bm{s}_{t})}\Bigg{[}\sum_{k=1}^{K}-r^{\mathcal{T}}(\bm{a}^{0}_{t})\cdot$
$\text{max}\Big{(}\rho_{k}(\theta,\theta_{\text{old}}),\text{clip}\left(\rho_{k}(\theta,\theta_{\text{old}}),1+\epsilon,1-\epsilon\right)\Big{)}\Bigg{]},$ (12)
@@
Fine-tuning the model solely depending on the reward is insufficient since the model may step too far, which can lead to performance collapse and instability during reward maximization. To address this problem, we introduce a Behavior-Clone (BC) regularization term during the fine-tuning process. Concretely, we aim to constrain our policy $\theta$ to closely match a target policy $\mu$, ensuring that $\theta$ does not deviate significantly from $\mu$ after policy updates. This constraint can be modeled using a negative log-likelihood (NLL) loss as:
@@
where $\epsilon(\bm{a}_{\mu}^{k},k)$ represents the ground-truth noise added to $\bm{a}_{\mu}^{k}$ at timestep $k$.
How to select the target policy? Intuitively, an ideal target policy is the optimal policy that generates samples $x^{*}$ satisfying $\mathcal{C}(\bm{x}^{*})\geq\mathcal{C}(\bm{x})$ for all possible $\bm{x}$, where $\mathcal{C}(\bm{x})$ represents a measure of the performance or quality of the sample, such as the accumulated reward for action sequences. Since $\mu$ is unknown during fine-tuning, we approximate it by sampling action sequences $\bm{a}$ that satisfy $\mathcal{C}(\bm{a})\approx\mathcal{C}(\bm{a}^{*})$. In practice, we denote $\bm{a}^{*}$ as the best actions from recent play experience, such as those that yielded the highest $n$ rewards or successfully completed the given task. We then sample $\bm{a}^{k}$ from these proficient actions obtained from online interaction, nearly equivalent to sampling from $\mu$ to regularize the fine-tuning process. We also remark that the BC regularizer is not the only way to incorporate regularization into Eq. (3.2). For example, a Kullback–Leibler (KL) divergence between fine-tuned and pre-trained models, or a diffusion pre-train loss can be employed to regularize the fine-tuning process, as shown in text-to-image and text-to-speech generation (Fan et al., 2024; Chen et al., 2024). However, we find these regularization may cause the pre-trained planner trap in sub-optimal regions, hindering performance improvement. We will further discuss them in experiments.
Combining Eq. (3.2) with Eq. (14), the loss function for reward fine-tuning in downstream tasks $\mathcal{T}\sim p(\mathcal{T})$ is expressed as follows:
@@
where $\lambda$ is a weight coefficient. The overall process of pre-training and fine-tuning using SODP is summarized in Alg. 1. Since our goal is to generate complete trajectories rather than individual segments, we utilize a trajectory-level buffer (Zheng et al., 2022) for estimating the target policy $\mu$. Further, to ensure the accuracy of the approximation, we generate several proficient trajectories using the pre-trained model at the beginning of each iteration.